OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square. Show your work for full credit.

OpenStudy (anonymous):

first ask yourself what two numbers that are added or subtracted can give you the sum or difference of 4 and when multiplied give you a product of 12

OpenStudy (anonymous):

To solve this by completing the square: 1. Put the constant (in this case, -12) on the other side: x^2 + 4x = 12 2. Take the coefficient of the x-term (in this case, 4), HALVE it, then SQUARE it: 4/2 = 2 2^2 = 4 3. Add that to BOTH sides: x^2 + 4x + 4 = 12 + 4 = 16 4. Simplify: x^2 + 4x + 4 = 16 (x + 2)^2 = 16 5. Take the square root of both sides, and put in a ± on the right side: (x+2)^2 = 16 x + 2 = ± 4 6. Simplify: x = -2 ± 4 7. Evaluate: x = -2 - 4 = -6 x = -2 + 4 = 2